OpenStudy (anonymous):

medal and fan what is 4 1/5 + 3 1/10=

OpenStudy (anonymous):

I CAN DO THIS...

OpenStudy (anonymous):

7 3/10

OpenStudy (anonymous):

or not...

OpenStudy (anonymous):

ok are you sure

OpenStudy (anonymous):

i smell math...

OpenStudy (anonymous):

can you guys help me with 4 more

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Dont just give the answer, that is against the code of conduct.

OpenStudy (anonymous):

he can

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

uhhh hold on

OpenStudy (anonymous):

bananas...

OpenStudy (anonymous):

what is 1 7/9 - 4/9=

OpenStudy (anonymous):

(banana x 3.14) divided by (monkeys + flying squirrels) to the third power.

OpenStudy (anonymous):

is it 1 11/9

OpenStudy (anonymous):

Solve your equation (for the first problem) 4 1/5 + 3 1/10= \[\frac{ 21}{ 5 }+3\frac{ 1 }{ 10}\] \[\frac{ 21 }{ 5 }+\] \[\frac{ 73 }{ 10 }=7\frac{ 3 }{ 10 }\]

OpenStudy (anonymous):

Sorry, its not bananas in the equation, its grapefruit.

OpenStudy (anonymous):

Solve: 1 7/9 - 4/9= \[\frac{ 16 }{9 }-\frac{ 4 }{ 9 }\] \[\frac{ 4 }{ 3 }\] \[1\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

\[1\frac{ 1 }{ 3}\] Is your answer.

OpenStudy (anonymous):

In decimal that is 1.333333

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Your welcome. Please open your next question in a new tab.

OpenStudy (anonymous):

ok